Suppose the demand and supply for good x are functions p=6.25-1/8Qd and p=1.75+1/10Qs what is the equilibrium price and quantity


1
Expert's answer
2021-05-31T14:48:34-0400

At equilibrium {quantity demands = quantity supplied }


That's "{Qd=Qs}"


"p=6.25-\frac{1}{8}Qd"


"-\frac{1}{8}Qd=p-6.25"


"Qd= -8p+50.....\u2026.................(1)"




"p= 1.75+\frac{1}{10}Qs"


"\frac{1}{10}Qs=p-1.75"


"Qs=10p-17.5...............................(2)"



"Qs=Qd"


"-8p+50=10p-17.5"


"18p=67.9"


"p(equilibrium" "price)= 3.75"


"Equilibrium" "quantity =10(3.75)-17.5=20"